OpenStudy (anonymous):

I just forgot how to solve these. ǀ2m - 8ǀ = 8 Can someone take me step to step?

OpenStudy (anonymous):

m could equal two answers

OpenStudy (anonymous):

since the absolute value could be 2m−8=8 or 2m−8=−8

OpenStudy (anonymous):

OH! Yea. I think I remember now. You have to make two problems right? ǀ2m - 8ǀ = 8 and ǀ2m - 8ǀ = -8